OpenStudy (anonymous):

A train leaves New York for Boston, 200 miles away, at 4:00 P.M. and averages 40 mph. Another train leaves Boston for New York on an adjacent set of tracks at 3:00 P.M. and averages 70 mph. At what time will the trains meet? (Round to the nearest minute.)

OpenStudy (anonymous):

the train which has left at 3:00 pm has already covered 70 miles so the distance left between 200-7=130

OpenStudy (anonymous):

since the trains are travelling in opposite direction their relative speed is 40+70 = 110 mph and the distance between them is now 130 miles

OpenStudy (anonymous):

now time = distance /speed

OpenStudy (anonymous):

the train which has left at 3:00 pm has already covered 70 miles so the distance left between 200-70=130